1. A series of dilute NaCl solutions are prepared starting with an initial stock solution of 0.100 M NaCl. Solution A is prepared by pipeting 10 mL of the stock solution into a 250-mL volumetric flask and diluting to volume. Solution B is prepared by pipeting 25 mL of solution A into a 100-mL volumetric flask and diluting to volume. Solution C is prepared by pipeting 20 mL of solution B into a 500-mL volumetric flask and diluting to volume. What is the molar concentration of NaCl in solutions A, B and C?

2. What is the molar concentration of NO3in a solution prepared by mixing 50.0 mL of 0.050 M KNO3 with 40.0 mL of 0.075 M NaNO3? What is pNO3 for the mixture?

3. To determine the concentration of ethanol in cognac a 5.00 mL sample of cognac is diluted to 0.500 L. Analysis of the diluted cognac gives an ethanol concentration of 0.0844 M. What is the molar concentration of ethanol in the undiluted cognac?
Answer #1

A = (0.100 M) × (10 mL) / (250 mL) = 0.0040 M

B = (0.0040 M) × (25 mL) / (100 mL) = 0.0010 M

C = (0.0010 M) × (20 mL) / (500 mL) = 4.0 × 10-5 M